# On the Robustness of Language Models for Tabular Question Answering

### v2

@@ Abstract
Large Language Models (LLMs), already shown to ace various text comprehension tasks have also remarkably been shown to tackle table comprehension tasks without specific training. While previous research has explored LLM capabilities with tabular dataset tasks, our study assesses the influence of in-context learning, model scale, instruction tuning, and domain biases on Tabular Question Answering (TQA). We evaluate the robustness of LLMs on Wikipedia-based WTQ, financial report-based TAT-QA, and scientific claims-based SCITAB, TQA datasets, focusing on their ability to interpret tabular data under various augmentations and perturbations robustly. Our findings indicate that instructions significantly enhance performance, with recent models exhibiting greater robustness over earlier versions. However, data contamination and practical reliability issues persist, especially with WTQ. We highlight the need for improved methodologies, including structure-aware self-attention mechanisms and better handling of domain-specific tabular data, to develop more reliable LLMs for table comprehension.
On the Robustness of Language Models for Tabular Question Answering
@@ 1 Introduction
LLMs, primarily trained on unstructured data, have performed fairly well with structured data like tabular datasets. Tasks that rely on tabular structures require the ability to process and understand data with structural awareness, even when the input is provided in a sequential linear format. Exploring such capability—particularly how LLMs apprehend and interpret structured information—can be effectively analyzed by studying robustness measures in table comprehension. This approach offers insights into the models’ ability to navigate and interpret the inherent relationships within structured datasets Borisov et al. (2023).
Figure 1: Example of different possible table augmentation methods where each row operation represents original table augmentation, transpose-based table augmentation, and value-based perturbation. The value-based perturbation only affects the cell content associated with the answer value of the given question.
Tabular Question Answering (TQA) is a popular task on tables, which involves answering a query based on the given structured data Herzig et al. (2020); Fang et al. (2024). TQA task highlights the emergent properties of LLMs, demonstrating their capability to comprehend structured data through in-context learning, instruction adherence, and multi-step reasoning Fang et al. (2024); Wei et al. (2022); Wang et al. (2024); Jiang et al. (2023); Dong et al. (2024); Liu et al. (2024). This capability can be quantitatively evaluated by measuring the robustness of the model’s performance on tabular datasets. Wang et al. (2022) recently showed the vulnerability of LLMs to structural augmentation of the input table and introduced a novel methodology to leverage structure-aware self-attention of an encoder-decoder model like T5 to more effectively manage table transformations, thereby improving the performance of natural language generation from tables. Similarly, Zhao et al. (2023) identified the vulnerabilities of recent TQA LLMs when faced with human-crafted adversarial perturbations. Liu et al. (2023) discusses inherent limitations within LLMs for tabular data interpretation, arguing that these models struggle with structural perturbations and advocate for an integrated approach that combines textual and symbolic reasoning. Chen (2023) find the performance of few-shot reasoning with LLMs for TQA show strong, complex reasoning abilities over table structures without table-specific pre-training.
Although these studies showcase a critical need for enhanced methodologies in handling tabular data across various applications, they have yet to fully explore how the robustness of performance with the inherent ability of LLMs varies with different factors. Notably, there has been little comparison of robustness as a function of (1) in-context learning, (2) the scale and instruction-tuning of LLMs, and (3) the biases inherent within existing evaluation benchmarks for robust table-to-text generation. These aspects are crucial, as the scale of LLMs can influence their ability to generalize from data, while instruction-tuning enhances in-context learning ability, affecting how well they adapt to table comprehension without explicit re-training. Additionally, biases in training data can skew the evaluation of their comprehension ability. Hence, understanding these dimensions is essential for developing more reliable and fair LLMs in processing tabular data Zhou et al. (2024). To evaluate these, Wikipedia-based datasets are the most popular benchmark datasets for TQA. We experiment with WTQ dataset Pasupat and Liang (2015), a Wikipedia-based tabular question-answering dataset, TAT-QA Zhu et al. (2021), a tabular question-answering dataset derived from financial reports, and SCITAB Lu et al. (2023) a scientific tabular question-answering dataset containing challenging claims that require compositional reasoning over tables. A general and specialized domain ensures reliable benchmarking of the robustness and generalizability of existing LLMs.
@@ 2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension, illustrated in Fig. 1.
2.1 Structural Perturbation (SP)
SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
2.2 Value Perturbation (ValP)
ValP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore these types of ValP:
Data Type Preserving Perturbation (DVP) [^1] Example of prompts and details of DVP dataset generation are present in the Appendix A, B: DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 6, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. We use such counterfactual entities to test the faithfulness of the LLM to the tabular data. 222 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’. Random Value Perturbation (RVP): RVP (an example is shown in Fig. 7) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
@@ 2.2 Value Perturbation (ValP)
3 Results
Figure 2: The average EM score across three different datasets for the different few shot comparisons for various perturbations with three different datasets.
3.1 Effects of ICL examples on TQA
Does instruction prompting assist LLM for better table comprehension for question-answering tasks? Fig. 11 compares the performance of various LLMs under different few-shot settings, measured by exact match (EM) scores. The plot illustrates the performances in 0, 1, 2, and 3-shot scenarios. Notably, there’s an increasing trend in EM with the number of shots, demonstrating that additional examples improve table comprehension abilities. Following the patterns observed in the WTQ dataset, the overall performance trends for both TAT-QA (Fig.13) and SCITAB (Fig.14) are generally consistent. However, within SCITAB, there are a few exceptions where smaller models outperform their larger counterparts. In general, we demonstrate complementary results to LLMs on unstructured tasks Wei et al. (2022) in the context of TQA. Fig. 2 and 12 focuses on the impact of various perturbations on model performance. Here, the models display a drop in EM across different manipulation operations, which indicates a sensitivity to changes in the data structure, underscoring a potential vulnerability in model robustness against input invariances and equivariances. The consistency in performance degradation across operations suggests that while models can handle the original task, their structural awareness ability still requires substantial enhancement for broader table comprehension.
@@ 3.2 Effects of the Model Type on TQA
Does newer models have better TQA abilities? Fig. 3 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Does instruction tuning help TQA? The heat maps in Fig. 4 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. Figure 4(right) distinctly indicate that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
How do performances vary with value perturbations?  [^3] More details are in Appendix Section F and G Table 1 and 4 show the performance of small and large LLMs, as defined on Table 3, when subjected to various types of value-based perturbation on WTQ using different evaluation metrics, Exact Match(EM), Variation percentage (VP) Yang et al. (2022); Zhou et al. (2024), and Exact Match Difference (Emd)Zhao et al. (2023); Zhou et al. (2024). Both model sizes experience a decline in exact match scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency offers resilience against total comprehension failure. VP increases significantly across all perturbations for both large and small models, indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios, reflecting the models’ vulnerability to these disruptions. The observed discrepancies in performance, particularly the deviation from the original model’s performance in scenarios like DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with manipulated table values.
ValP Large Model Small Model
Operation EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - -
NT 0.05 0.36 -0.31 0.02 0.25 -0.23
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23
Table 1: The average EM, VP, and Emd of small and large LLMs on Value Perturbations(ValP). The classification of large and small models is defined in Table 3.
@@ 3.3 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 outlines the reliance of LLMs on structured tabular data, particularly evident in the NT and NVP operations. When no table is provided, both type of models show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5of queries correctly, indicating a potential bias towards Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the original structured format without the relevant data provides enough contextual cues for the models to answer, highlighting model biases.
@@
How do LLMs perform on (TAT-QA) and (SCITAB), which is a specialized dataset that requires stronger reasoning abilities? Both large and small LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset(Fig 4(middle)) and SCITAB dataset(Fig 4(right)). For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in the domain-specific dataset. Instruction-tuned models, demonstrate enhanced effectiveness, reflecting the advantages of model-specific optimization for harder tabular comprehension tasks. From Table 4, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals a critical insight: while models start with higher EM on WTQ but suffer larger Emd and higher VP, indicating higher sensitivity to perturbations, their relatively smaller Emd and lower VP on both TAT-QA and SCITAB reflect greater robustness—despite SCITAB ’s overall lower baseline EM.. Particularly the sharp contrast in the NT performance, where on TAT-QA and SCITAB dataset, models have approximately 0% accuracy, a distinguishable contrast from the WTQ performance. This suggests that performance metrics on WTQ might be inflated due to biases favoring familiarity with Wikipedia’s structured data, whereas TAT-QA and SCITAB, a more specialized rarer domain, is more challenging. This highlights the need for better benchmark datasets to assess LLMs’ robust table comprehension abilities accurately.
@@ 4 Conclusion
Our study explores the table comprehension abilities of LLMs through the lens of various invariant and equivariant perturbations. We find that instruction tuning and ICL examples enhance the table comprehension ability of LLMs. Newer models, such as Llama3, are more robust compared to their predecessors. Additionally, we uncover domain biases for evaluating tabular data using Wikipedia-based datasets. Despite the decent performance of LLMs on TQA, they still fall short of the reliability required for several practical applications.
@@ 5 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to including a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain specific dataset and Wikipedia-based dataset. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structural aware or fine-tuned models for tabular dataset. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks.

### v3

# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

@@ Abstract
Large Language Models (LLMs), already shown to ace various text comprehension tasks, have also remarkably been shown to tackle table comprehension tasks without specific training. Building on earlier studies of LLMs for tabular tasks, we probe how in-context learning (ICL), model scale, instruction tuning, and domain bias affect Tabular QA (TQA) robustness by testing LLMs, under diverse augmentations and perturbations, on diverse domains: Wikipedia-based WTQ, financial TAT-QA, and scientific SCITAB. Although instruction tuning and larger, newer LLMs deliver stronger, more robust TQA performance, data contamination and reliability issues, especially on WTQ, remain unresolved. Through an in-depth attention analysis, we reveal a strong correlation between perturbation-induced shifts in attention dispersion and the drops in performance, with sensitivity peaking in the model’s middle layers. We highlight the need for improved interpretable methodologies to develop more reliable LLMs for table comprehension.
Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis
@@
footnotetext: * These authors jointly supervised this work.
@@ 1 Introduction
LLMs, despite being primarily trained on unstructured text, have demonstrated notable capabilities in structured-data tasks such as Tabular Question Answering (TQA). TQA requires models to interpret data presented in tables, demanding strong structural reasoning. Assessing how LLMs navigate structured comprehension challenges can provide valuable insights into their robustness and reasoning capabilities Borisov et al. (2023); Fang et al. (2024).
Recent studies highlight the importance of robustness evaluations in understanding LLM behavior on structured tasks Zhou et al. (2024). Specifically, perturbations in tabular structure and content significantly impact model performance, highlighting vulnerabilities that limit practical reliability Wang et al. (2022); Zhao et al. (2023). Furthermore, Liu et al. (2023) argue that LLMs inherently struggle with structural manipulations, advocating for an integrated approach combining symbolic reasoning to enhance model robustness.
Figure 1: (a) An example of the different possible table augmentation methods.(b) Scatter plot between change in attention entropy and EM score across perturbation types for Llama-8B-Instruct on WTQ.
Although these studies identify vulnerabilities and suggest potential improvements, there remains limited understanding of how internal model mechanisms respond to perturbations in structured data. Attention mechanisms form the core of transformer-based LLMs, governing how models distribute focus across input elements during processing Clark et al. (2019). Prior work analyzing attention patterns in natural language tasks revealed that specific layers and attention heads critically influence model performance and robustness Zhao et al. (2024); Barbero et al. (2025). However, detailed attention-level analyses for structured tasks like TQA remain scarce.
To address this gap, we systematically investigate LLM robustness for TQA tasks across multiple dimensions, including in-context learning, model scale, instruction tuning, domain biases, value perturbations, and model size. Our experiments evaluate different perturbation types, highlighting how value-based alterations influence reasoning fidelity and faithfulness (Table 1). We also compare performance across diverse datasets to assess domain biases and generalizability.
Extending beyond surface-level performance metrics, we quantify changes in attention entropy across different attention heads and layers under various perturbations, analyzing how these changes correlate with performance degradation. With this we aim to uncover which attention heads are most sensitive to perturbations, informing targeted improvements for future models.
Our study leverages diverse datasets: Wikipedia-based WTQ Pasupat and Liang (2015), financial report-based TAT-QA Zhu et al. (2021), and scientific claims-based SCITAB Lu et al. (2023). Through these diverse domains, our findings give insights into the sensitivity and reliability of LLMs on TQA, highlighting the broader challenge of understanding how LLMs reason over structured data.
@@ 2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension, illustrated in Fig. 1. Structural Perturbation (SP): SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
Value Perturbation (ValP): ValP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore these types of ValP:
Data Type Preserving Perturbation (DVP) [^1] Example of prompts and details of DVP dataset generation are present in the Appendix A, B: DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 5, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. These counterfactual entities test the faithfulness of the LLM to the tabular data. 222 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’.Random Value Perturbation (RVP): RVP (an example is shown in Fig. 6) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
@@
3 Evaluation Performance
Operation WTQ Dataset TATQA Dataset SCITAB Dataset
EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - - 0.33 - - 0.20 - - 0.098 - - 0.132 - -
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11 - - - - - - - - - - - -
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15 - - - - - - - - - - - -
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23 - - - - - - - - - - - -
Table 1: The average EM, VP, and Emd of small and large LLMs on all perturbation operations. Large and Small LLMs are defined on Table 3.
@@ 3.1 Effects of the Model Type on TQA
How do performances vary with value perturbations?  [^3] More details are in Appendices F and G Table 1 shows the performance of various LLMs, as defined on Table 3, when subjected to different value-based perturbation on WTQ using different evaluation metrics: Exact Match(EM), Variation percentage (VP) Yang et al. (2022); Zhou et al. (2024), and Exact Match Difference (EMD)Zhao et al. (2023); Zhou et al. (2024). We see LLMs experience a decline in EM scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency helps. VP increases significantly across all perturbations indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios. The observed discrepancies in performance, particularly pronounced in DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with perturbed tables.
@@ 3.2 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 shows that when no table is provided, LLMs show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5of queries correctly, indicating a potential bias in Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the models are biased by the contextual structured format cue even in the absence of relevant data.
@@
How do out-of-box LLMs perform on specialized domains: TAT-QA and SCITAB ? As shown in Table 1, LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset and SCITAB dataset. For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in niche domains, but instruction-tuned models are better here. From Table 1, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals an interesting accuracy-robustness tradeoff: while models with higher EM on WTQ suffer larger EMD and higher VP, indicating higher sensitivity to perturbations, their relatively smaller EMD and lower VP on both TAT-QA and SCITAB reflect greater robustness, despite SCITAB ’s overall lower baseline EM. Also notable is the sharp contrast in the NT performance, where on TAT-QA and SCITAB dataset, models have approximately 0% accuracy, a contrast from the WTQ performance. This suggests that performance on WTQ might be inflated due to biases favoring familiarity with Wikipedia, compared to niche domains like TAT-QA and SCITAB. This highlights the need for better benchmark design for tabular understanding. We also observe trends consistent with prior work: performance improves with few-shot prompting, instruction-tuned models generally outperform their base counterparts, and larger models tend to exhibit stronger TQA Wei et al. (2022); Fang et al. (2024). Additional results and full breakdowns across datasets, model sizes, and fine-tuning configurations are included in Appendix G.
4 Attention Analysis
Figure 2: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset.
Understanding how structural perturbations in tabular inputs affect a model is critical for assessing robustness and diagnosing potential failure modes in TQA. As attention mechanisms regulate how language models allocate focus across table elements, examining their sensitivity to perturbations provides insights into representational stability. Here, we analyze this sensitivity by quantifying how perturbation-induced changes in attention dispersion correlate with performance degradation.
4.1 Effect of Perturbations on Attention Maps
We examine how varying severities of structural perturbations influence attention weights using the Llama3-8B-Instruct model on the WTQ dataset. For each attention head across all layers, we measure the change in attention entropy (detailed in Appendix H) between the original and perturbed table inputs. In parallel, we compute the corresponding change in Exact Match (EM) scores, capturing the performance impact of each perturbation.
Attention entropy captures the distribution uniformity of attention across tokens; higher entropy indicates diffuse attention, while lower entropy signifies focused attention. Thus, significant entropy changes reflect considerable shifts in the model’s internal attentional focus due to perturbations.
Our analysis (Figure 1) reveals a significant positive correlation (Spearman $\rho=0.90$, $p=0.037$) between changes in attention entropy and EM degradation across perturbation types. This shows that more severe perturbations significantly disrupt the model’s attention distribution, consequently diminishing its ability for TQA.
These findings highlight the sensitivity of attention mechanisms to structural integrity within tabular inputs. Perturbations affecting relational semantics induce greater attention dispersion and misalignment, directly impairing model accuracy. Thus, attention dispersion is a crucial indicator linking input perturbations with performance.
4.2 Layer-Wise Analysis of Attention Head Sensitivity to Perturbations
Figure 3: Layer-wise Spearman correlation between aggregated attention entropy change (averaged over heads) and EM difference.
While averaging attention metrics across heads provides a general understanding of model behavior, prior research emphasizes that individual attention heads and specific layers, particularly the middle layers, significantly contribute to encoding task-relevant information Barbero et al. (2025); Zhao et al. (2024). To investigate this at a finer granularity, we perform a layer-wise analysis of how perturbation-induced changes in attention entropy correlate with EM score degradation.
Specifically, we use the Llama3-8B-Instruct model on the WTQ dataset to compute the Spearman correlation between entropy changes (original vs. perturbed inputs) and corresponding EM differences for each attention head across all layers. Figure 2 demonstrates that correlations peak predominantly in the middle layers and remain consistently elevated through these central layers. This indicates that perturbation-induced shifts in attention distribution within middle layers are strongly predictive of performance degradation. Additional experiments are included within the Appendix H.
To further substantiate this pattern, we aggregate entropy changes across all heads within each layer and analyze the correlation with EM differences (Figure 3). This aggregated analysis reinforces the observation of heightened sensitivity in mid-layer attention mechanisms.Collectively, these results highlight that perturbation sensitivity is not uniformly distributed across model layers. Instead, the middle layers, which typically engage in critical representational refinement, are particularly vulnerable to attention instability and associated performance degradation.
@@ 5 Conclusion
Our study offers a comprehensive robustness analysis of LLMs under various perturbations for TQA. While larger, instruction-tuned models show improved performance, they remain highly sensitive to structural and value-based disruption. We also uncover domain biases, where models perform well on WTQ without tables yet struggle on more specialized datasets like TAT-QA and SCITAB. We show that shifts in attention entropy, particularly in middle layers, correlate with performance degradation. This work highlights the importance of interpretability and robustness of LLMs for TQA.
@@ 6 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to include a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain-specific datasets and Wikipedia-based datasets. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structurally aware or fine-tuned models for tabular datasets. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for the question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks. Moreover, we only conduct a case study of attention analysis with small models because of computation cost.